OpenStudy (anonymous):

sin(pi/2-x)=sin(pi/2+x)

OpenStudy (anonymous):

yes, true

OpenStudy (anonymous):

you can use the "addition angle" formula to prove it if you like

OpenStudy (anonymous):

\[\sin(y-x)=\sin(y)\cos(x)-\cos(y)\sin(x)\] put \(y=\frac{\pi}{2}\)